# Memory-less and Backscatter-less Tunnel Diode Harmonic Signatures for RFID

### v1

@@ Abstract
Tunnel diodes have traditionally been researched for extending backscatter read-ranges for ultra-high-frequency (UHF) radio-frequency identification (RFID) tags as reflection amplifiers. This paper explores the natural harmonics that arise from biasing these diodes within their negative differential resistance regions and with no interrogating signal from a transmitting source, such as an RFID reader, to injection-lock these diodes. These harmonics are characterized for five tunnel diode boards, made with the same components and with each board’s fundamental frequencies measuring at above -15 dBm at a biasing voltage of 200 mV when measured over-the-cable. The occurrence of these harmonics creates unique harmonic signatures for each board and demonstrates possible harmonic RFID applications that can help RFID readers discover and even identify RFID tags with backscatter-less and memory-less IDs generated by tunnel diodes.
@@ I Introduction
Tunnel diodes are heavily-doped semiconductor devices that exhibit negative resistance at low biasing voltages, typically starting below 300 mV. They are capable of producing reflection coefficient magnitudes greater than unity at low turn-on voltages. Thus, they are used to make the tunneling reflection amplifiers that have commonly been studied for extending the read-ranges of passive radio-frequency identification (RFID) tags [1, 2, 3, 4]. Tunnel diodes can and should be used to address issues, other than path loss, that affect the backscatter link quality between tag and reader. Eid et al. [3] has investigated leveraging tunnel diodes’ low turn-on voltages for energy-harvesting rectification and negative resistance for compensating circuit losses in building an oscillator. Saetia and Durgin [5] have exploited tunneling reflection amplifiers to create multi-bit symbols for amplitude-phase shift keying (APSK). Self-jamming, identifying-tags in cluttered environments, etc. are some other issues in RFID systems that tunnel diodes can be used to help address. Specifically, this paper aims to highlight the characterization of tunnel diodes’ natural resonance harmonics and investigate using those harmonics for memory-less identification of tags, offering an alternative method to generate tags’ IDs and sense the presence of tags in an area, all backscatter-less.
@@ I-A The Harmonic RFID Concept
Researchers have been exploring the use of frequency harmonics to better identify tags for both chipless and chipped tag designs. Chipless RFID tags typically use different resonators to create a unique frequency response signature. As noted by Barbot et al. [6], these chipless tags are linear and time-invariant, making their read-ranges limited compared to traditional passive ultra-high frequency (UHF) RFID tags that are non-linear. Being non-linear allows traditional UHF tags to backscatter information on frequencies outside the carrier frequency of the interrogating signal. This use of non-carrier frequencies improves these tags’ detection since the reader’s receiver can distinguish the tags’ weaker backscattered signals from the directly interfering signal from the reader’s transmitter and the other reflected signals from objects in the environment. Barbot et al. [6] noted that the inclusion of a nonlinear element to a chipless tag allows the tag to be nonlinear. The use of schottky diodes to allow chipless tags to backscatter at the second-order harmonic of the fundamental/interrogating signal has been demonstrated in [7] and [8].
@@
Both aspects have been investigated to improve identification of these tags. Barbot [9] explored having the RFID reader detect passive tags after interrogating their nonlinear chips and looking for the tags’ backscattered frequency components that lie away from the interrogation carrier signal’s frequency. Piumwardane et al. [10] varied duty cycles of the toggling of a tag’s load modulator to create different harmonic/frequency signatures to assign tags with unique IDs.
Tunnel diodes, like any non-linear device, produces harmonics. In 1963, Guttmann [11] suggested the use of tunnel diodes as harmonic generators due to their nonlinearity and negative resistance/conductance. More recently, Gumber et al. [12] demonstrated the ability for tunnel diodes to backscatter on a fundamental and second harmonic to help resolve the RFID reader’s self-jamming problem through frequency multiplexing the transmit and receive operations. Inspired by that work, this research aims to further investigate the idea of creating a harmonic tunneling tag for identification. For this work, compared to [12], emphasis was placed on measuring the tunnel diode’s frequency response or natural resonance when only given a DC-bias; thus:
@@
    No injection-locking with an intentional external RF interrogating signal is used.
@@
    No matching network was used to guide where the tunneling reflection amplifier has high reflection gains or resonances.
@@ II The Tunnel Diode Oscillator Model
In this work, the tunnel diodes are laid on boards with a RF choke, a DC smoothing capacitor, and microstrip lines. Only a DC biasing voltage is needed to make these boards resonant, similar to how oscillators convert DC power to an AC output signal [13]. Thus, to the best of the authors’ knowledge, the circuit model and reasoning for the frequency generation by this work’s tunnel diode boards can be best explained by the established one-port oscillator model. Negative resistance is often used to create the one-port oscillator model, as depicted in Fig. 2.
@@
$X_{res}+R_{res}(I,j\omega)=0$ (3)
Since $R_{res}$ is a positive resistance value, then $R_{N}(I,j\omega)$ must be less than zero, negative resistance. The impedance values from the added $Z_{N}$ cancels out the impedance values of $Z_{res}$, allowing for a non-decaying current flow. Importantly, due to the impedances’ dependence on frequency and current, specific frequencies and voltages/currents allow for conditions (2) and (3) to be met to ensure continuous oscillation at $\omega=\omega_{o}$, the frequency of steady oscillation. For steady-state, single-frequency stability, the Kurokawa’s condition, fully derived in [13], helps ensure the oscillator maintains a stable resonance at the steady-state frequency, despite perturbation effects from applied frequency and current changes to the circuit. Because tunnel diodes, like any diodes, are non-linear devices, harmonics of the fundamental resonating frequency should appear on the devices’ measured output spectrum, unless filtered out.
Tunnel diodes, in their current-voltage curves, have negative-differential regions (NDRs) that can provide the negative resistance in this one-port oscillator model at specific voltages. [14, 15, 3] demonstrated the construction of tunnel diode oscillators. Even hobbyists tinkering with old Russian tunnel diodes can build simple tunnel diode oscillators. For example, [16] showcased a tunnel diode oscillator built by simply looping capacitors and a wire inductor around a tunnel diode’s leads, making a harmonic-generating and tunable oscillator that is dependent on the biasing voltages. This same tunable and harmonic-generating behavior will be observed in this work.
@@ III Board Design
The schematic in Fig. 3 shows the layout of the tunnel diode boards, and Table I lists their parameters. The parameters of the AI101E tunnel diode used can be found at [17]. Each tunnel diode board, as pictured in Fig. 4, was originally designed to be reflection amplifiers for [5]. Because of the circuit setup of the DC-biasing and tunnel diode portions of the board, these reflection amplifier boards also can operate as oscillators. The dual-functionality of tunnel diodes to serve as the basis of reflection amplifiers and oscillators had been noted by Eid et al. in [3]. It is the intention of this work for these tunnel diode boards to serve as both reflection amplifier backscatter modulators and harmonics-generators.
Interestingly, all boards were designed to be of the same specifications in Table I. Any differences in these boards come from the manufacturing and fabrication processes of their components and how they were even assembled on the board. These slight differences will cause different identifiable responses, as will be shown in the later sections.
@@ IV-A Cabled Measurement Setup
Figure 5: Measurement setup.
Shown in Fig. 5, an Agilent N9000A Signal analyzer, with a DC block, is connected to the tunnel diode board. A DC power supply biased the tunnel diode at various voltages. The measurements setup parameters for the signal analyzer and the DC power supply are listed in Table II.
@@
Biasing Voltage Sweep 0-300 mV (1 mV increments)
@@
Notably, for all these measurements, no interrogation signal is sent to the boards. These boards are simply biased by the power source, and it was predicted that the voltages needed to bias the diodes in the negative differential region (NDR) will lead to the appearance of harmonic content in their measured spectrum. This behavior is evident as will be shown in the next section that showcases these boards’ measurements at 200 mV, specifically.
IV-B Harmonic Behavior at 200 mV
@@
At 200 mV, the captured spectrum of the five AI101E reflection amplifier boards are plotted in Fig. 6. These output spectrum measurements, along with the ones in Fig. 7, did not ignore the combined cable and DC block insertion loss, which was at most 0.2 dB. Shown by this figure and the fundamental frequencies listed in Table III, each board had varying fundamental frequencies and different number of harmonics within the measured 3 GHz span. Note, the fundamental frequency is defined as the highest-amplitude, non-DC frequency content. Each board’s fundamental frequency appears to be at least above -15 dBm.
@@
IV-C Spectrum-Voltage Signatures
Figure 7: Performing a frequency and voltage sweep, a spectrum color map for each board is created. The pattern of warm-colored stripes show the harmonic signature for each board across a range of voltages. The green dotted circles highlight the stable harmonics, while the black dash-dot loops highlight the jittery harmonic response as the tunnel diodes enter their NDRs.
To investigate what happens when different voltages are applied to the tunneling reflection amplifiers, the signal analyzer’s frequency sweeps for each biasing voltage sweep value are plotted on a color-map in Fig. 7. As seen in that figure, there are lighter-colored ”stripes” for where the tunnel diodes are resonating. From 0 volts to about 150-180 mV, depending on the boards, the diodes are not in the negative differential region; thus, the boards are not resonating. But then as the boards are biased closer to the biasing range of the NDRs, the spectral response becomes jittery before stabilizing.
A close-up of each board’s fundamental frequency demonstrates that the different biasing voltages can tune the fundamental frequency location by biasing them with different values within their diodes’ negative differential resistance regions. Board 1 has the widest tunable range of about 60 MHz, while Board 3 has the narrowest tunable range of about 1 MHz.
Figure 8: Shifting the fundamental frequency by applying different biasing voltages within their NDRs.
@@ V-A Wireless Measurement Setup
Two wide-band, high-gain horn antennas, which are rated to operate from 1-18 GHz and provide over 7 dBi of gain past 1 GHz [18], were spaced 64 inches or 1.63 meters apart in an anechoic chamber, shown in Fig. 9. One antenna was connected to the tunnel diode board, the transmitter (TX). The other antenna was connected to either a spectrum analyzer or an oscilloscope, the receiver (RX). Due to the limited physical space of the chamber, 1.63 meters is the furthest the two antennas were spaced from each other. Shown in Fig. 10, the tunnel diode boards, the spectrum analyzer, and oscilloscope were all placed outside the anechoic chamber during testing.
Two sets of measurements were taken in the frequency and time domains with, respectively, an Agilent E4446A spectrum analyzer and a Tektronix MSO64 oscilloscope that has 8 GHz bandwidth channels and areal-time sampling rate of 25 GS/s. For the wireless spectrum measurements, the following parameters in Table IV were used to set-up the spectrum analyzer and the DC power supply. For the time-domain/transient measurements taken with the oscilloscope, the horizontal span and trigger were adjusted manually to best capture the waveform.
@@
TABLE IV: Wireless measurement setup parameters.
@@ V-B Wireless-Measured Spectrum Results
Fig. 11 shows the wirelessly measured harmonic signatures of the tunnel diode boards at 200 mV. Boards 1, 2, and 4 from the wireless measurements have notable harmonic signatures at 200 mV. Boards 3 and 5 do demonstrate frequency-producing behavior at 200 mV, but not distinct harmonics as their spectrum content is spread/clustered over several frequencies. This behavior is possibly due to impedance mismatch between those specific boards and the antenna and cable. Fig. 12 demonstrates the harmonic signature behavior of the boards over several biasing voltages. Because of the path loss from these wireless measurements, the receive strength of the harmonic signatures was not as strong as the measured over-the-cable signatures captured in Fig. 7. To aid with visualization, the green dotted loops in Fig. 12 highlights the stable harmonics of the boards, while the black dash-dot loops mark the jittery-wide-band spectrum behavior, demonstrated for example by Boards 3 and 5 in Fig. 11. As expected, Boards 1 and 2 have stable, distinct harmonic signatures when biased in their negative differential resistance regions. The jittery, wide-band behavior demonstrated by Boards 3 and 5, even in their NDRs, could be because of impedance mismatch with the antenna and cable as the biasing voltage on the board changes the board’s impedance.
It should be noted that because the antennas have below 0 dBi of gain at frequencies below about 700 MHz [18], sub-1 GHz fundamental components will not appear as strong as they did in the over-the-cable measurement results. Table V shows the predicted fundamental frequency locations based on these wireless measurements and the location of the visible second harmonics. Also, the harmonics locations for the boards measured wireless occurred at different locations than those that were measured over-the-cable due to the change in testing setup with attaching the antenna and longer cable with different impedances. Such as shown in Fig. 8 with the over-the-cable measurements, Fig. 13 demonstrates the tunability of the tunnel diode boards with the wireless measurement results. It should be noted the center frequencies in that figure are of the strongest frequency component in the captured spectrums of Fig. 11, so they are not always the fundamental frequencies as were used in Fig. 8.

### v2

@@ Abstract
[…]
@@ I Introduction
Tunnel diodes are heavily-doped semiconductor devices that exhibit negative resistance at low biasing voltages, typically starting below 300 mV. They are capable of producing reflection coefficient magnitudes greater than unity at low turn-on voltages. Thus, they are used to make the tunneling reflection amplifiers that have commonly been studied for extending the read-ranges of passive radio-frequency identification (RFID) tags [1, 2, 3, 4]. Tunnel diodes can and should be used to address issues, other than path loss, that affect the backscatter link quality between tag and reader. For example, Eid et al. [3] investigated leveraging tunnel diodes’ low turn-on voltages and negative resistance, respectively, for rectification and compensating resistance losses in an oscillator design. Saetia and Durgin [5] have exploited tunneling reflection amplifiers to create multi-bit symbols for amplitude-phase shift keying (APSK). Self-jamming, identifying-tags in cluttered environments, etc. are some other issues in RFID systems that tunnel diodes can be used to help address. Specifically, this paper aims to highlight the characterization of tunnel diodes’ natural resonance harmonics and investigate using those harmonics for memory-less identification of tags, offering an alternative method to generate tags’ IDs and sense the presence of tags in an area, all backscatter-less.
@@ I-A The Harmonic RFID Concept
Researchers have been exploring the use of frequency harmonics to better identify both chipless and chipped tag designs. Chipless RFID tags typically use different resonators to create a unique frequency response signature. As noted by Barbot et al. [6], these chipless tags are linear and time-invariant, making their read-ranges limited compared to traditional passive ultra-high frequency (UHF) RFID tags that are non-linear. Being non-linear allows traditional UHF tags to generate harmonic and inter-modulated signals and, thus, backscatter information on frequencies outside the carrier frequency of the interrogating signal. This use of non-carrier frequencies improves these tags’ detection since the reader’s receiver can distinguish the tags’ weaker backscattered signals from the directly interfering signal from the reader’s transmitter and the other reflected signals from objects in the environment that match the carrier frequency. Barbot et al. [6] noted that the inclusion of a nonlinear element to a chipless tag allows the tag to be nonlinear. The use of schottky diodes to allow chipless tags to backscatter at the second-order harmonic of the fundamental/interrogating signal has been demonstrated in [7] and [8].
@@
Both aspects have been investigated to improve identification of these tags. Barbot [9] explored having the RFID reader detect passive tags after interrogating their nonlinear chips and looking for the tags’ backscattered frequency components that are generated around the interrogation carrier signal’s frequency. Piumwardane et al. [10] varied duty cycles of the toggling of a tag’s load modulator to create different harmonic/frequency signatures to assign tags with unique IDs.
Tunnel diodes, like any non-linear device, produces harmonics. In 1963, Guttmann [11] suggested the use of tunnel diodes as harmonic generators due to their nonlinearity and negative resistance/conductance. More recently, Gumber et al. [12] demonstrated the ability for tunnel diodes to backscatter on a fundamental and second harmonic to help resolve the RFID reader’s self-jamming problem through frequency multiplexing the transmit and receive operations. Inspired by that work, this research aims to further investigate the idea of creating a harmonic tunneling tag for identification. In contrast to [12], emphasis was placed on measuring the tunnel diode’s frequency response or natural resonance when only given a DC-bias; thus:
@@
    No injection-locking or any intentional external RF interrogating signal was used on the designed tunnel diode boards. These boards simply served as oscillators with no backscattering involved.
@@
    No matching networks were used to determine the frequencies these tunnel diodes would have high reflection gains and resonance strengths.
@@ II The Tunnel Diode Oscillator Model
In this work, the tunnel diodes are soldered onto printed circuit boards with a RF choke, a DC smoothing capacitor, and microstrip lines. Only a DC biasing voltage is needed to make these boards resonant, similar to how oscillators convert DC power to an AC output signal [13]. Thus, to the best of the authors’ knowledge, the circuit model and reasoning for the frequency generation by this work’s tunnel diode boards can be best explained by the well-established one-port oscillator model. Negative resistance is often used to create the one-port oscillator model, as depicted in Fig. 2.
@@
$X_{res}(j\omega)+X_{N}(I,j\omega)=0$ (3)
Since $R_{res}$ is a positive resistance value, then $R_{N}(I,j\omega)$ must be less than zero, negative resistance. The impedance values from the added $Z_{N}$ cancels out the impedance values of $Z_{res}$, allowing for a non-decaying current flow. Importantly, due to the impedances’ dependence on frequency and current, specific frequencies and voltages/currents allow for conditions (2) and (3) to be met to ensure continuous oscillation. For steady-state and single-frequency stability where $\omega=\omega_{o}$, the frequency of steady oscillation, the Kurokawa’s condition, fully derived in [13], helps ensure the oscillator maintains a stable resonance at the steady-state frequency, despite perturbation or changes to the circuit’s frequency or current. Because tunnel diodes, like any diodes, are non-linear devices, harmonics of the fundamental resonating frequency should appear on the devices’ measured output spectrum, unless filtered out.
Tunnel diodes, in their current-voltage curves, have negative differential resistance (NDR) regions that can provide the negative resistance in this one-port oscillator model at specific voltages. [14, 3, 15, 16, 17] demonstrated the construction of tunnel diode oscillators. Specifically, Eid et al. [3] and Varshney et al. [17] both demonstrated the capability to use a tunnel diode as both a low-power reflection amplifier and oscillator for low-power wireless communications. Additionally, Varshney and Corneo [17] used a simple tunnel diode oscillator that is injection-locked by a nearby backscattered signal from a tag to relay a stronger response to a receiver. For $\geq$100 GHz and even THz communications, research has been conducted to build high-frequency, resonant-tunneling-diode (RTD) oscillators that can output tens of micro-watts [18] to greater than half a milli-watt[19, 20] of RF power. Even amateur hobbyists have used old Russian tunnel diodes to build simple tunnel diode oscillators. For example, [21] showcased a tunnel diode oscillator built by simply looping capacitors and a wire inductor around a tunnel diode’s leads, making a harmonic-generating and tunable oscillator that is dependent on the biasing voltages. This same tunable and harmonic-generating behavior will be observed in this work.
@@ III Board Design
The schematic in Fig. 3 shows the layout of the tunnel diode boards, and Table I lists their parameters. The parameters of the AI101E tunnel diode used can be found at [22]. Each tunnel diode board, as pictured in Fig. 4, was originally designed to be reflection amplifiers for [5]. Because of the circuit setup of the DC-biasing and tunnel diode portions of the board, these reflection amplifier boards also can operate as oscillators. The dual-functionality of tunnel diodes to serve as the basis of reflection amplifiers and oscillators had been noted by Eid et al. [3] and Varshney et al. [17]. It is the intention of this work for these tunnel diode boards to serve as both reflection amplifier backscatter modulators and harmonics-generators.
Interestingly, all boards were designed to be of the same specifications in Table I. Any differences in these boards came from the manufacturing and fabrication processes of their components and how they were even assembled on the board. These slight differences will cause different identifiable responses, as will be shown in the later sections.
@@ IV-A Cabled Measurement Setup
Figure 5: Cabled measurements setup.
Shown in Fig. 5, an Agilent N9000A Signal analyzer, with a DC block, is connected to the tunnel diode board under test. A DC power supply biased the tunnel diode at various voltages. The measurements setup parameters for the signal analyzer and the DC power supply are listed in Table II.
@@
Biasing Voltage Sweep 3-300 mV (1 mV increments)
@@
Notably, for all these measurements, no interrogation or injection-locking signal is sent to the boards. These boards were simply biased by the DC power supply, and it was predicted that the voltages needed to bias the diodes in the negative differential resistance (NDR) regions will lead to the appearance of harmonic content in their measured spectra. This behavior is evident as will be shown in the next section that showcases these boards’ measurements at 200 mV, specifically.
IV-B DC Voltage, Current, and Power Consumption
Figure 6: Top figure plots the measured DC current drawn by the tunnel diodes in relation to the measured applied DC biasing voltages. Bottom figure plots the measured DC power consumed by the tunnel diodes.
Fig. 6(a) shows the measured DC current as each board was swept through different biasing voltages. Likewise, Fig. 6(b) demonstrates the calculated DC power consumption for each board from multiplying the measured DC current and biasing voltage values. Device-dependent, the tunnel diodes started to enter their negative differential resistance regions at around 100 mV, then fully into their NDR regions when $\geq$150 mV of biasing voltage was applied, evident in Fig. 6(a) by the sharp negative slope in the current-voltage (IV) curves. For Fig. 6(b), in the NDR regions, the curves plateaued after a steep rise prior to entering the NDR regions.
IV-C Harmonic Behavior at 200 mV
@@
At 200 mV, the captured spectrum of the five AI101E reflection amplifier boards are plotted in Fig. 7. Note, these output spectrum measurements, along with the ones in Fig. 8, did not exclude the combined cable and DC block insertion losses, which was at most 0.2 dB and negligible. Shown by this figure and the fundamental frequencies listed in Table III, each board had varying fundamental frequencies and different number of harmonics within the measured 3 GHz span. Note, the fundamental frequency is defined as the highest-amplitude, non-DC frequency content. Each board’s fundamental frequency appeared to be at least above -15 dBm.
@@
IV-D Spectrum-Voltage Signatures
Figure 8: Performing a frequency and voltage sweep, a spectrum color map for each board is created. The pattern of warm-colored stripes show the harmonic signature for each board across a range of voltages. The green dotted circles highlight the stable harmonics, while the black dash-dot loops highlight the jittery harmonic response as the tunnel diodes enter their NDR regions.
To investigate what happens when different voltages are applied to the tunnel diode boards, the measured spectrum from the signal analyzer for each biasing voltage sweep value are plotted as a color-map in Fig. 8. As seen in that figure, there are lighter-colored ”stripes” for where the tunnel diodes are resonating. Up to 150-180 mV, depending on the specific board, the diodes were not biased into their negative differential resistance regions; thus, the boards were not resonating. But then as the boards are biased closer to the biasing range of their NDR regions, the spectral response became jittery before stabilizing.
A close-up of each board’s fundamental frequency in Fig. 9 demonstrates that the different biasing voltages can tune the fundamental frequency location by biasing them with different values within their diodes’ negative differential resistance regions. Board 1 has the widest tunable range of about 60 MHz, while Board 3 has the narrowest tunable range of about 1 MHz. Varshney and Corneo [16] noted similar frequency tuning for the AI101D tunnel diodes based on input biasing voltage changes.
Figure 9: Shifting the fundamental frequency by applying different biasing voltages within the tunnel diodes’ NDR regions.
IV-E DC-to-RF Efficiency
(a) $\eta_{DC-RF}$ in relation to biasing voltage.
(b) $\eta_{DC-RF}$ in relation to DC power consumed.
[…]
$\eta_{DC-RF}=\frac{P_{peak,out}}{P_{DC}}$ (4)
[…]
1 30.04 664.6 295.9
2 23.53 597.7 273.1
3 26.53 758.5 279.9
4 29.58 779.6 287.4
5 17.63 782.3 252.9
TABLE IV: DC-RF Efficiency Values
[…]
@@ V-A Wireless Measurement Setup
Two wide-band and high-gain horn antennas were spaced 64 inches or 1.63 meters apart in an anechoic chamber, shown in Fig. 11. These antennas are specifically rated to be 50 ohms, operate from 1-18 GHz, and provide over 7 dBi of gain past 1 GHz [24]. One antenna was connected to the tunnel diode board, the transmitter (TX). The other antenna was connected to either a spectrum analyzer or an oscilloscope, the receiver (RX). Due to the limited physical space of the chamber, 1.63 meters is the furthest the two antennas were spaced from each other. Shown in Fig. 12, the tunnel diode boards, the spectrum analyzer, and oscilloscope were all placed outside the anechoic chamber during testing. The TX cable connecting the tunnel diode board to the TX antenna had at most 3.68 dB of insertion loss, and the RX cable connecting the RX antenna had at most 2.24 dB of insertion loss. Note, these losses were not accounted for in the following captured spectra and transient measurements.
Two sets of measurements were taken in the frequency and time domains with, respectively, an Agilent E4446A spectrum analyzer and a Tektronix MSO64 oscilloscope that has 8 GHz bandwidth channels and a real-time sampling rate of 25 GS/s. For the wireless spectrum measurements, the following parameters in Table V were used to set-up the spectrum analyzer and the DC power supply. For the time-domain/transient measurements taken with the oscilloscope, the horizontal span and trigger were adjusted manually to best capture the waveform.
@@
TABLE V: Wireless measurement setup parameters.
@@ V-B Wireless-Measured Spectrum Results
Fig. 13 shows the wireless measured harmonic signatures of the tunnel diode boards at 200 mV. Boards 1, 2, and 4 from the wireless measurements have notable harmonic signatures at 200 mV. Boards 3 and 5 do demonstrate frequency-producing behavior at 200 mV, but not distinct harmonics as their spectrum content is spread/clustered over several frequencies. The differences in the measured over-the-cable and wireless results can be possibly due to impedance mismatch between those specific boards and the antenna and cable and/or the antenna’s impedance modifying the overall resonant behavior of the boards when the antenna was loaded onto the boards. Fig. 14 demonstrates the harmonic signature behavior of the boards over several biasing voltages. Because of path loss and cable losses, the receive strength of the harmonic signatures was not as strong as the measured over-the-cable signatures captured in Fig. 8. To aid with visualization, the green dotted loops in Fig. 14 highlights the stable harmonics of the boards, while the black dash-dot loops mark the jittery-wide-band spectrum behavior, demonstrated for example by Boards 3 and 5 in Fig. 13. As expected, Boards 1 and 2 have stable, distinct harmonic signatures when biased in their negative differential resistance regions. The jittery, wide-band behavior demonstrated by Boards 3 and 5, even in their NDR regions, could be because of impedance mismatch with the antenna and cable as the biasing voltage on the board changes the board’s impedance.
It should be noted that because the antennas have below 0 dBi of gain at frequencies below about 700 MHz [24], sub-1 GHz fundamental components will not appear as strong as they did in the over-the-cable measurement results. Table VI shows the predicted fundamental frequency locations based on these wireless measurements and the location of the visible second harmonics. Also, the harmonics locations for the boards measured wireless occurred at different locations than those that were measured over-the-cable due to the change in testing setup with attaching the antenna and longer cable with different impedances. Such as shown in Fig. 9 with the over-the-cable measurements, Fig. 15 demonstrates the tunability of the tunnel diode boards with the wireless measurement results. Visually inspecting Fig. 15, Board 1 demonstrated the best tunable range at about 5 MHz. It should be noted the center frequencies in that figure are of the strongest frequency component in the captured spectra of Fig. 13, so they are not always the fundamental frequencies as were used in Fig. 9.